the sum of four consecutive odd integers is -72. write an equation to model this situation, and find the values of the four integers
Answer by Alan3354(69443)

the sum of four consecutive odd integers is -72. write an equation to model this situation, and find the values of the four integers
==========
n + n+2 + n+4 + n+6 = -72
4n + 12 = -72
-----
or
============
-72/4 = -18, the center
There are 2 below and 2 above it.
--> -15, -17, -19, -21